# GRAND for Gaussian Intersymbol Interference Channels

### 10 Mar 2026 v2

@@ III-C Properties of Sequence Reliability
If a subset $\mathcal{W}\subseteq\left\{1,\cdots,N\right\}$ satisfies $\mathcal{W}=\left\{a,a+1,\cdots,a+\lvert\mathcal{W}\rvert-1\right\},a\in\mathcal{N}^{+}$, then $\mathcal{W}$ is called non-decomposable and we have
@@ IV GRAND-ISI For Higher-Order ISI Channels
  • The sets $\mathcal{I}_{i}$ form a partitioning of $\mathcal{I}(\underline{e}(m))$, i.e.,
@@ IV-A Properties of Sequence Reliability
For first-order ISI channels in Section III, $\mathcal{W}$ is called non-decomposable if it satisfies $\mathcal{W}=\left\{a,a+1,\cdots,a+\lvert\mathcal{W}\rvert-1\right\}$ for some $a\in\mathcal{N}_{+}$. The definition of non-decomposability for higher-order ISI channels remains the same. Furthermore, there arises the partially-decomposable case, as described below:
@@
We obtain $\underline{x}^{\star}=1001$ and the sequence reliabilities are:
@@ V-C Complexity Analysis
In this subsection, we provide an analysis of algorithm complexity. The procedure involved in ORBGRAND-ISI can be broken down into three main steps: (1) acquiring the hard detection sequence, (2) calculating the sequence reliability for each error burst, and (3) generating and testing the burst patterns. Compared with ORBGRAND-ISI, CDF-ORBGRAND-ISI and 2line-ORBGRAND-ISI only require additional table-lookup according to Fig. 1, with rather marginal additional cost.
@@
For ORBGRAND-AI, as described in [7], its procedure can also be divided into the same three steps. Therefore, we can compare its complexity with ours. Taking $\mathrm{E_{b}/N_{0}}=4$dB in Fig. 2a as an example, the results are summarized in Table II, where Step 1 refers to the complexity of the Viterbi algorithm (assuming $L\leq\text{block length}$), Step 2 denotes the number of sequence reliabilities to be computed (or reliabilities in ORBGRAND-AI), and Step 3 reports the average number of queries (each entry is given in the form of candidate/valid, corresponding to the average numbers of queries for candidate EPs and valid EPs, respectively). We observe that ORBGRAND-ISI, 2line-CDF-ORBGRAND-ISI, and CDF-ORBGRAND-ISI incur similar computational complexity. Compared to ORBGRAND-AI12, the number of computed sequence reliabilities is reduced by about a factor of five, and the corresponding number of queries is also reduced accordingly.
@@ VII-A Proof of Theorem 1
$=$ $\sum_{t=1}^{\zeta(\mathcal{I}(\underline{e}(m)))}\sum_{i=\min(\mathcal{I}_{t}(\underline{e}(m)))}^{\text{R}_{1}(\mathcal{I}_{t}(\underline{e}(m)))}\ln\frac{P_{\mathsf{Y}_{i}|\underline{\mathsf{X}}_{i-1:i}}(y_{i}|\underline{x}^{*}_{i-1:i})}{P_{\mathsf{Y}_{i}|\underline{\mathsf{X}}_{i-1:i}}(y_{i}|[f_{\mathcal{I}_{t}}(\underline{x}^{*})]_{i-1:i})}$

### 11 Mar 2026 v3

@@ III-C Properties of Sequence Reliability
If a subset $\mathcal{W}\subseteq\left\{1,\cdots,N\right\}$ satisfies $\mathcal{W}=\left\{a,a+1,\cdots,a+\lvert\mathcal{W}\rvert-1\right\},a\in\mathbb{N}_{+}$, then $\mathcal{W}$ is called non-decomposable and we have
@@ IV GRAND-ISI For Higher-Order ISI Channels
  • The sets $\mathcal{I}_{i}$ form a partitioning of $\mathcal{I}(\underline{e})$, i.e.,
@@ IV-A Properties of Sequence Reliability
For first-order ISI channels in Section III, $\mathcal{W}$ is called non-decomposable if it satisfies $\mathcal{W}=\left\{a,a+1,\cdots,a+\lvert\mathcal{W}\rvert-1\right\}$ for some $a\in\mathbb{N}_{+}$. The definition of non-decomposability for higher-order ISI channels remains the same. Furthermore, there arises the partially-decomposable case, as described below:
@@
We obtain $\underline{x}^{\star}=0100$ and the sequence reliabilities are:
@@ V-C Complexity Analysis
In this subsection, we provide an analysis of algorithm complexity. The procedure involved in ORBGRAND-ISI can be broken down into three main steps: (1) acquiring the hard detection sequence, (2) calculating the sequence reliability for each error burst, and (3) generating and testing the EPs. Compared with ORBGRAND-ISI, CDF-ORBGRAND-ISI and 2line-ORBGRAND-ISI only require additional table-lookup according to Fig. 1, with rather marginal additional cost.
@@
For ORBGRAND-AI, as described in [7], its procedure can also be divided into the same three steps. Therefore, we can compare its complexity with ours. Taking $\mathrm{E_{b}/N_{0}}=4$dB in Fig. 2a as an example, the results are summarized in Table II, where Step 1 refers to the complexity of the Viterbi algorithm (assuming $L\leq\text{block length}$), Step 2 denotes the number of sequence reliabilities to be computed (or reliabilities in ORBGRAND-AI), and Step 3 reports the average number of queries (each entry is given in the form of candidate/valid, corresponding to the average numbers of queries for candidate EPs and valid EPs, respectively). We observe that ORBGRAND-ISI, 2line-CDF-ORBGRAND-ISI, and CDF-ORBGRAND-ISI incur similar computational complexity. Compared to ORBGRAND-AI12, the number of computed reliabilities is reduced by about a factor of five, and the corresponding number of queries is also reduced accordingly.
@@ VII-A Proof of Theorem 1
$=$ $\sum_{t=1}^{\zeta(\mathcal{I}(\underline{e}(m)))}\sum_{i=\min(\mathcal{I}_{t}(\underline{e}(m)))}^{\text{R}_{1}(\mathcal{I}_{t}(\underline{e}(m)))}\ln\frac{P_{\mathsf{Y}_{i}|\underline{\mathsf{X}}_{i-1:i}}(y_{i}|\underline{x}^{*}_{i-1:i})}{P_{\mathsf{Y}_{i}|\underline{\mathsf{X}}_{i-1:i}}(y_{i}|[f_{\mathcal{I}_{t}(\underline{e}(m))}(\underline{x}^{*})]_{i-1:i})}$